# Binarization-Aware Adjuster: Bridging Continuous Optimization and Binary Inference in Edge Detection

### v1

@@ Abstract
Image edge detection (ED) faces a fundamental mismatch between training and inference: models are trained using continuous-valued outputs but evaluated using binary predictions. This misalignment, caused by the non-differentiability of binarization, weakens the link between learning objectives and actual task performance. In this paper, we propose a theoretical method to design a Binarization-Aware Adjuster (BAA), which explicitly incorporates binarization behavior into gradient-based optimization. At the core of BAA is a novel loss adjustment mechanism based on a Distance Weight Function (DWF), which reweights pixel-wise contributions according to their correctness and proximity to the decision boundary. This emphasizes decision-critical regions while down-weighting less influential ones. We also introduce a self-adaptive procedure to estimate the optimal binarization threshold for BAA, further aligning training dynamics with inference behavior. Extensive experiments across various architectures and datasets demonstrate the effectiveness of our approach. Beyond ED, BAA offers a generalizable strategy for bridging the gap between continuous optimization and discrete evaluation in structured prediction tasks.
I Introduction
Image edge detection (ED) is a fundamental task in computer vision, serving as a critical component for higher-level applications such as segmentation, inpainting, and object recognition [1, 2, 3]. While ED has advanced significantly—from early classical methods [4, 5] to modern deep learning-based approaches [6, 7, 8, 9, 10], a core theoretical inconsistency remains underexplored.
This inconsistency lies in the gap between the continuous-valued outputs used for training and the binary predictions required at inference. Ground-truth edge maps are binary, indicating the presence or absence of edges at each pixel. However, to enable gradient-based optimization, models are trained to produce continuous values in $[0,1]$, typically using loss functions such as binary cross-entropy (BCE). Because binarization is discontinuity, it is excluded from the training loop and applied only post hoc during evaluation. This disconnect can lead to suboptimal performance: a model that minimizes the training loss may not yield the most accurate binary predictions after thresholding.
In practice, a fixed threshold is used to convert predicted probabilities into a binary edge map. Predictions near this threshold are inherently uncertain and more prone to misclassification, making them especially critical to the final binary outcome. In contrast, confidently correct predictions, namely those far from the threshold and on the correct side, have limited impact on the binary result, but they often dominate the loss due to their ease of optimization. As a result, standard continuous-value loss functions allocate excessive learning capacity to uninformative regions while neglecting the more decision-critical predictions near the decision threshold. This imbalance impairs the effectiveness of training and may ultimately degrade binary performance.
While prior work has improved ED performance through architectural innovations [6, 7, 10] and alternative loss functions [11, 12, 13], few have addressed the theoretical misalignment between continuous-valued training objectives and binary evaluation metrics. The binarization step remains disconnected from the optimization process, leaving a fundamental gap in the learning framework.
To address this, we propose a theoretical method to design a binarization-aware adjuster (BAA) that explicitly incorporates binarization behavior into the training process. Central to our method is a novel threshold-aware loss adjustment mechanism that dynamically modulates each pixel’s contribution to the loss based on its proximity to the decision threshold and correctness. Predictions near the threshold or those misclassified receive higher weights, while confidently correct predictions are down-weighted. This principled adjustment redirects the learning signal toward decision-critical regions, aligning optimization objectives with the binary nature of the task.
Our main contributions are as follows:
  • We identify a core mismatch in ED: the disconnect between continuous-valued supervision in model training and binary-valued final output in practical usage, and propose the BAA method to resolve it.
  • We introduce a principled method for constructing BAA based on Distance Weight Function (DWF), and demonstrate how it generalizes and improves upon loss formulations.
  • We validate our approach through extensive experiments across multiple benchmark datasets and architectures, demonstrating the effectiveness of the proposed method.
II Previous Works
This section reviews the landscape of ED research across datasets, model architectures, loss functions, and evaluation protocols, with an emphasis on foundational developments and persisting challenges.
II-A Datasets
Early ED research relied heavily on datasets designed for related tasks such as contour detection and image segmentation. Notable benchmarks like BSDS300 and BSDS500 [14], featuring multi-annotator contour labels, remain widely used. Other commonly adopted datasets include MDBD [15] for boundary detection, NYUD [16] for semantic segmentation, and general-purpose datasets such as PASCAL VOC [17], COCO [18], SceneParse150 [19], and Cityscapes [20]. However, these were not originally designed for ED.
As the field matured, specialized ED datasets emerged. BIPED and its improved version BIPED2 [21] introduced high-resolution single-edge annotations. BRIND [6] extended the annotations of BSDS500 with semantically labeled edges, including reflectance, illumination, normal, and depth. UDED [22] offered fewer samples but prioritized annotation quality and consistency, enabling more reliable evaluation.
@@ II-A Datasets
Nevertheless, annotation quality remains a key bottleneck. Human-labeled edge maps often suffer from noise, inconsistency, and semantic ambiguity. Although several studies have proposed more robust training strategies and refined labeling protocols [23, 24, 25], annotation noise continues to challenge both model generalization and fair benchmarking.
II-B Model Architectures
The evolution of ED models can be divided into three phases: classical heuristics, early learning-based methods, and deep learning-driven approaches.
Classical methods such as the Sobel operator [4] and Canny detector [5] used hand-crafted gradient filters to capture intensity discontinuities. While computationally efficient, these methods lack adaptability to semantic structures and struggle in complex visual environments.
Early learning-based models incorporated statistical learning using manually crafted features like texture descriptors [26], sketch tokens [27], or histogram-based cues [28], paired with classifiers such as structured forests [29] or logistic regression [30]. However, these approaches were fundamentally limited by their reliance on hand-engineered features and had limited generalization across diverse scenes.
The advent of deep learning, particularly convolutional neural networks (CNNs), opened a new path for ED. HED [11] introduced deep supervision across network layers from multiple resolutions, while RCF [31] and BDCN [32] refined the multi-scale fusion techniques. Models such as PiDiNet [33] and DexiNet [34] focused on designing new structures for improving ED prediction. Beyond CNNs, transformer-based methods [7, 35] and diffusion-based approaches [8, 9] explore global context modeling and have also been employed in ED-related tasks, though their advantages in the ED task itself remain under study. Recently, the Extractor-Selector (E-S) paradigm [10] introduced a two-stage framework combining feature extraction with a learnable selection module, offering a new perspective on the challenge of discriminative feature selection.
II-C Loss Functions
Loss function design plays a critical role in ED. The Weighted Binary Cross-Entropy (WBCE) loss remains the most widely adopted approach, emphasizing edge pixels to address the low recall rate of the BCE loss. Despite its effectiveness in obtaining a high recall rate, WBCE often requires non-maximum suppression (NMS) at inference time to obtain clean, thin edge maps, resulting in a trade-off between quantitative score and prediction quality.
To address this, various alternative loss functions have been proposed. Dice loss [36] emphasizes region-level overlap, while tracing loss [37] targets the proportion of edges within edge-adjacent areas. Ranking-based approaches, such as RankED [12] and AP-loss [38], aim to enhance precision by optimizing pixel-level ranking. More recently, SWBCE [13] accounts for human uncertainty by rebalancing the contributions of edge and non-edge regions.
While these advanced losses provide perceptual or structural improvements, they often introduce new challenges such as instability, reduced precision, or sensitivity to hyperparameters, hindering their widespread use. A unified loss that aligns training supervision with binary evaluation, while maintaining stability and generalization, remains an open problem.
II-D Evaluation Metrics
Standard evaluation metrics for ED were established by [26] and remain widely adopted. The two most common metrics are:
@@ II-D Evaluation Metrics
  • ODS (Optimal Dataset Score): Best F1-score computed using a single optimal binarization threshold applied across the entire dataset.
@@
  • OIS (Optimal Image Score): Mean of best F1-scores computed per image using image-specific optimal thresholds.
These metrics are computed using an error-tolerance matching strategy to allow minor spatial misalignments. However, the error-tolerance radius is often dataset-specific and can influence results and comparability. Early works did not unify the error-tolerance benchmark among datasets until recent studies advocated for a fixed 1-pixel tolerance to ensure stricter, more consistent evaluation [25].
III Methodology
III-A Motivation and Insight
@@
A key challenge in training ED models is the mismatch between training objectives and evaluation protocols. During training, ED models generate continuous-valued outputs in the range $[0,1]$, typically optimized using pixel-wise loss functions such as WBCE. In contrast, evaluation metrics are computed on binarized edge maps, obtained by applying a fixed threshold to these outputs. Since the binarization operation is discontinuous, it cannot be directly incorporated into the training process. This disconnect creates a fundamental issue: the loss function does not explicitly align with the discrete classification decisions that ultimately determine evaluation performance.
@@
This gap introduces a critical inefficiency in optimization. Pixels whose predicted values lie on the correct side and far from the decision threshold, namely those correct pixels with high confidence, tend to dominate the gradient signal since they are easy to optimize. In contrast, pixels with predictions near the threshold, which are more ambiguous and critical to classification, contribute less to the gradients since they are more difficult to optimize. Consequently, the optimizer, driven by gradient magnitude, inadvertently prioritizes ’easy’ but less effective pixels over ’ambiguous’ but crucial ones, due to the nature of the optimization that tends to optimize the less-challenging pixels with higher priority. This misallocation of optimization effort can prevent the optimal parameter be learned, resulting in suboptimal learning and degraded final performance.
@@
For example, Table I reports optimal binarization thresholds for various ED models on multiple datasets under WBCE loss, which typically fall between 0.6 and 0.8. Suppose the binarization threshold is 0.7, then reducing the predicted value of a non-edge pixel from 0.50 to 0.01 has no effect on the final binary edge map, but it significantly reduces the loss. Moreover, such confident correct predictions are often easy to optimize. Indeed, it is because they are easy to optimize that they become the correctly predicted ones with high confidence. Therefore, training naturally prioritizes them for their optimization ease, even though further improvement on them yields no benefit for the final decision. In contrast, increasing the predicted value of an edge pixel from 0.69 to 0.71 can be more difficult, which is the reason why its prediction is near the decision threshold. Therefore, though critical for correct classification, it contributes less to the loss and thus is more easily overlooked.
@@
TABLE I: Optimal binarization thresholds under WBCE training: Evaluation uses 1-pixel error tolerance without NMS. Experimental settings follow the ones in Section IV-A. Models are retrained on the pre-training set and evaluated on the validation set.
@@
III-B Adjusted Loss Function Formulation
Although designing the BAA is non-trivial, the idea of using an adjuster is conceptually simple. Let us first address the simpler task, formally defining the adjusted loss function in this subsection, and leave the more challenging part, designing the BAA, to the next.
@@ III-B Adjusted Loss Function Formulation
$L_{Adj,\delta}(Pred,Gt)=\sum_{i}(Adj(P_{i},G_{i})+\delta)L(P_{i},G_{i})$ (1)
@@
where $i$ runs over all pixels, $L$ is a selected baseline pixel-wise loss function such as WBCE [^1] The same symbol $L$ is employed to denote both the loss function on a single pixel ($L(P_{i},G_{i})$) and on the whole image ($L(Pred,Gt)$)., $Pred$ and $Gt$ represent the model prediction and the ground truth, respectively, $P_{i}$ and $G_{i}$ represent the model prediction and the ground truth of the $i$-th pixel, respectively, and $\delta\geq 0$ is a pre-defined hyperparameter, directly adding loss $L$ with intensity $\delta$ to the adjusted loss to stabilize training. The function $Adj$ is called an Adjuster, where $Adj(P_{i},G_{i})\in[0,1]$ assigns dynamic weights, emphasizing the training signal according to the pixel’s contribution to the final binary decision.
@@
Despite the simple idea, the problem lies in the design of the adjuster. A naive choice of $Adj(P_{i},G_{i})$ is the binary mask $\hat{Adj}_{thr}(P_{i},G_{i})$, defined as:
$\hat{Adj}_{thr}(P_{i},G_{i})=\left\{\begin{matrix}&0&If\ (P_{i}-thr)(G_{i}-thr)\geq 0\\ &1\ &Otherwise\end{matrix}\right.$ (2)
@@
where $thr$ is the binarization threshold. Hence, $L_{Adj,\delta}$ satisfying (providing $\delta=0$):
  • If a pixel is correctly classified, then the loss on it is 0.
  • If a pixel is incorrectly classified, then the loss on it is positive.
@@
This adjuster emphasizes only misclassified pixels. However, the discontinuity of its corresponding adjusted loss function renders it unsuitable for gradient-based optimization. Therefore, a smoother, differentiable formulation is necessary.
III-C Design of the Binarization-Aware Adjuster

### v2

# Binarization-Aware Adjuster: A Theoretical Framework for Bridging Continuous Optimization and Discrete Inference with Application to Edge Detection

@@ Abstract
In machine learning, discrete decision-making tasks exhibit a fundamental inconsistency between training and inference: models are optimized using continuous-valued outputs, yet evaluated through discrete predictions. This discrepancy arises from the non-differentiability of discretization operations, weakening the alignment between optimization objectives and practical decision outcomes. To address this, we present a theoretical framework for constructing a Binarization-Aware Adjuster (BAA) that integrates binarization behavior directly into gradient-based learning. Central to the approach is a Distance Weight Function (DWF) that dynamically modulates pixel-wise loss contributions based on prediction correctness and proximity to the decision boundary, thereby emphasizing decision-critical regions while de-emphasizing confidently correct samples. Furthermore, a self-adaptive threshold estimation procedure is introduced to better match optimization dynamics with inference conditions. As one of its applications, we implement experiments on the edge detection (ED) task, which also demonstrate the effectiveness of the proposed method experimentally. Beyond binary decision tasks and ED, the proposed framework provides a general strategy for aligning continuous optimization with discrete evaluation and can be extended to multi-valued decision processes in broader structured prediction problems.
keywords:
Adjusted Loss Function, Binarization-Aware Adjuster, Distance Weight Function, Continuous-Discrete Alignment, Edge Detection
[…]
1 Introduction
Many real-world vision tasks, such as semantic segmentation, image classification, object recognition, and edge detection (ED) [1, 2, 3, 4, 5], ultimately require discrete decisions (e.g., binary or categorical outputs). However, learning methods, such as deep neural networks, constrained by the need for differentiable computation, are typically optimized using continuous-valued loss functions within gradient-based learning frameworks [6, 7, 8, 9, 10]. As non-differentiable inference operations, such as thresholding, quantization, or category assignment, are excluded from the optimization process, a fundamental gap arises between continuous-valued optimization during training and the discrete inference deployed at test time. This gap may lead to suboptimal performance, as optimization does not always align with decision-critical behavior.
This mismatch is particularly evident in pixel-level binary prediction tasks. For example, edge detection (ED) aims to generate binary edge maps but is typically trained using continuous-valued loss functions such as binary cross-entropy (BCE). Since binarization is non-differentiable, it is excluded from the optimization loop and applied only post hoc during evaluation. As a result, model updates tend to prioritize pixels that are easy to optimize but have limited influence on the binary output, whereas pixels near the decision boundary, which are more critical for final accuracy, receive insufficient emphasis due to their lower weights and higher optimization difficulty.
Several studies have attempted to mitigate the inconsistency between decision behavior and training objectives, typically by changing losses that rebalance decision classes, such as weighted BCE [11] or alternative ones [12, 13]. However, to the best of our knowledge, few existing methods explicitly integrate binarization (or discretization) behavior into gradient-based optimization, leaving the fundamental optimization–inference misalignment unresolved.
To address this issue, we propose a theoretically grounded framework named Binarization-Aware Adjuster (BAA), which explicitly embeds binarization behavior into the optimization process. The core idea is a threshold-aware loss adjustment mechanism based on the proposed Distance Weight Function (DWF), which dynamically reweights pixel-level loss according to prediction correctness and proximity to the decision threshold. This enables the model to allocate greater learning capacity to decision-critical regions while down-weighting confidently correct predictions that exert limited influence on binary outputs. Additionally, a self-adaptive threshold estimation strategy is introduced to align the adjuster with inference behavior. Since the BAA is continuous and differentiable, it can be seamlessly incorporated into existing gradient-based optimization pipelines, allowing discrete inference behavior to influence learning directly.
The proposed framework is designed for binary decision tasks and is applied to the ED task to demonstrate its effectiveness. ED is selected as a representative implementation scenario due to its foundational role in low-level vision and its strong dependence on precise decision boundaries. Furthermore, many structured prediction tasks can be viewed as extensions or generalizations of the ED paradigm. Binary decision tasks are selected since they can display our method most conveniently and clearly. Nevertheless, the proposed BAA mechanism is conceptually extensible to multi-value discrete decision tasks and applicable to a broader range of tasks involving discrete decisions, such as segmentation, classification, and object recognition.
The main contributions of this work are summarized as follows:
  • 1.
    We identify a fundamental mismatch between continuous optimization and discrete inference in model training and propose the Binarization-Aware Adjuster (BAA) framework to bridge this gap.
  • 2.
    We develop a principled formulation based on the Distance Weight Function (DWF), enabling binarization-aware loss adjustment while retaining full differentiability.
  • 3.
    We apply the proposed framework to the edge detection task and implement extensive experiments on substantial benchmark datasets across multiple architectures.
2 Previous Works
A large number of learning-based systems perform discrete decision-making while being optimized using continuous-valued loss functions. This discrepancy has motivated studies on improving the alignment between continuous optimization during training and discrete evaluation at inference. In this work, edge detection (ED) is adopted as an illustrative validation scenario due to its strong dependency on binary decisions and its relevance to many structured prediction tasks. In the following, we briefly review related studies with ED as a representative example.
2.1 Datasets
Historically, early vision research commonly relied on general-purpose datasets. For instance, BSDS300/500 [14] have been used in contour detection, segmentation, and ED. Common datasets such as PASCAL VOC [15], COCO [16], and NYUD [17] have been widely adopted across segmentation, detection, and image classification tasks. With increasing task specialization, dataset selection became more focused, e.g., PASCAL VOC for object detection, NYUD for segmentation, and BSDS500 for contour-based analysis.
For ED specifically, dedicated datasets were developed to provide more accurate edge annotations. BIPED/BIPED2 [18] offers edge annotations for 250 high-resolution campus images, BRIND [6] extended BSDS500 with reannotated edge labels, and UDED [19] provides carefully selected high-quality samples.
@@ 2.1 Datasets
Despite these efforts, annotation quality remains a major challenge. Human-labeled edge maps often suffer from inconsistency, ambiguity, and noise. Although robust training strategies and refined labeling protocols have been proposed [20, 21, 22], annotation noise still affects both generalization ability and fair benchmarking.
2.2 Model Architectures
Deep learning architectures have achieved broad success in computer vision [23], natural language processing [24], and decision systems [25]. Major architectural paradigms, including convolutional neural networks (CNNs), transformer models, and diffusion-based methods, have demonstrated significant cross-domain applicability.
Using ED as an example, the pioneering work HED [11] validated the effectiveness of multi-scale deep supervision using CNNs, outperforming traditional approaches [4, 5] and early feature-based learning methods [26, 27, 28, 29, 30]. Subsequent works, such as RCF [31], BDCN [32], PiDiNet [33], and DexiNet [34], further optimized fusion strategies and model structures. Transformer-based [7, 35] and diffusion-based [8, 9] designs have also been explored. More recently, the Extractor–Selector framework [10] demonstrated improvements through refined feature selection. Among these approaches, CNN-based models remain the most widely adopted due to their robustness and maturity.
2.3 Loss Functions
Loss function design is critical in model optimization. Binary cross-entropy (BCE) was initially the standard for decision tasks. However, BCE assigns equal importance to positive and negative samples, which is unsuitable for imbalanced data. To address this, Weighted BCE (WBCE) is widely adopted and still the most standard one currently. Other alternatives include Dice loss [36] and tracing loss [37] for geometry-aware optimization, AP-loss [38] and RankED [12] based on ranking formulation, and SWBCE [13] that models uncertainty in manual annotations.
Although these methods improve sample weighting, they generally do not explicitly model the final inference rule, leaving the discretization step separate from the optimization process. Consequently, the fundamental continuous–discrete gap remains unresolved. This limitation motivates the development of our BAA framework.
2.4 Evaluation Metrics
[…]
$F\beta=\frac{(1+\beta^{2})\times\text{Precision}\times\text{Recall}}{\beta^{2}\times\text{Precision}+\text{Recall}},$ (1)
where
$\text{Precision}=\frac{TP}{TP+FP},\quad\text{Recall}=\frac{TP}{TP+FN}.$ (2)
Here, TP denotes true positives, FP false positives, and FN false negatives, with $\beta$ typically set to 1. The most frequently used metrics include:
  • 1.
@@ 2.4 Evaluation Metrics
    ODS (Optimal Dataset Score): Best F1 using a single threshold for the entire dataset.
  • 2.
@@
    OIS (Optimal Image Score): Mean of each image’s best F1 using image-specific thresholds.
For ED, the commonly used metric computation was proposed in [26] with error tolerance to compensate for minor spatial deviations. Recent studies [22] recommend stricter and standardized tolerance settings for more reliable comparison.
3 Methodology: General Theory
This section presents the theoretical framework of our method. Here, the design for binary decision tasks is proposed for illustrative convenience and simplicity. Note that the discussions can be easily extended to multi-value discrete decision tasks.
3.1 Motivation and Insight
@@
A key challenge in training binary decision models is the mismatch between training objectives and evaluation protocols. During training, models generate continuous-valued outputs, typically normalized to [0,1], and are optimized using pixel-wise loss functions such as WBCE. In contrast, evaluation metrics are computed on binarized maps, obtained by applying a fixed threshold to these outputs. Since the binarization operation is discontinuous, it cannot be directly incorporated into the training process. This disconnect creates a fundamental issue: the loss function does not explicitly align with the discrete classification decisions that ultimately determine evaluation performance.
@@
This gap introduces a critical inefficiency in optimization. Pixels whose predicted values lie on the correct side and far from the decision threshold, namely those correct pixels with high confidence, tend to dominate the gradient signal since they are easy to optimize. In contrast, pixels with predictions near the threshold, which are more ambiguous and critical to classification, contribute less to the gradients since they are more difficult to optimize. Consequently, the optimizer, driven by gradient magnitude, inadvertently prioritizes ’easy’ but less effective pixels over ’ambiguous’ but crucial ones, due to the nature of the optimization that tends to optimize the less-challenging pixels with higher priority. This misallocation of optimization effort can prevent the optimal parameter from being learned, resulting in suboptimal learning and degraded final performance.
@@
For example, Table 1 reports optimal binarization thresholds for various ED models on multiple datasets under WBCE loss, which typically fall between 0.6 and 0.8. Suppose the binarization threshold is 0.7, then reducing the predicted value of a non-edge pixel from 0.50 to 0.01 has no effect on the final binary edge map, but it significantly reduces the loss. Moreover, such confident-correct predictions are often easy to optimize, which is the reason that they become the correctly predicted ones with high confidence. Therefore, training naturally prioritizes them for their optimization ease since they contribute more to loss, even though further improvement on them yields no benefit for the final decision. In contrast, increasing the predicted value of an edge pixel from 0.69 to 0.71 can be more difficult, which is the reason why its prediction is near the decision threshold. Therefore, though critical for correct classification, it contributes less to the loss and thus is more easily overlooked.
@@
Table 1: Optimal binarization thresholds under WBCE training for ED models: Evaluation uses 1-pixel error tolerance without NMS. Experimental settings follow the ones in Section 4.1. Models are retrained on the pre-training set and evaluated on the validation set.
@@
3.2 Adjusted Loss Function Formulation
Although designing the BAA is non-trivial, the idea of using it is conceptually simple. Let us first address the simpler task, formally defining the adjusted loss function in this subsection, and leave the more challenging part, designing the BAA, to the next.
@@ 3.2 Adjusted Loss Function Formulation
$L_{Adj,\delta}(Pred,Gt)=\sum_{i}(Adj(Pred_{i},Gt_{i})+\delta)L(Pred_{i},Gt_{i})$ (3)
@@
where $i$ runs over all pixels, $L$ is a selected baseline pixel-wise loss function such as WBCE [^1] The same symbol $L$ is employed to denote both the loss function on a single element such as a pixel ($L(Pred_{i},Gt_{i})$) and on the whole object such an images ($L(Pred,Gt)$)., PredPred and $Gt$ represent the model prediction and the ground truth, respectively, $Pred_{i}$ and $Gt_{i}$ represent the model prediction and the ground truth of the $i$-th element, respectively, and $\delta\geq 0$ is a pre-defined hyperparameter, directly adding loss $L$ with intensity $\delta$ to the loss to stabilize training. The function $Adj$ is called an Adjuster, where $Adj(Pred_{i},Gt_{i})\in[0,1]$ assigns dynamic weights, emphasizing the training signal according to the pixel’s contribution to the final binary decision.
@@
Despite the simple idea, the problem lies in the design of the adjuster. A naive choice of $Adj(Pred_{i},Gt_{i})$ is the binary mask $\hat{Adj}_{thr}(Pred_{i},Gt_{i})$, defined as:
$\hat{Adj}_{thr}(Pred_{i},Gt_{i})=\left\{\begin{matrix}&0&If\ (Pred_{i}-thr)(Gt_{i}-thr)\geq 0\\ &1\ &Otherwise\end{matrix}\right$. (4)
@@
where thrthr is the binarization threshold. Hence, $L_{Adj,\delta}$ satisfying (providing $\delta=0$):
  • 1.
    If an element is correctly classified, then the loss on it is 0.
  • 2.
    If an element is incorrectly classified, then the loss on it is positive.
@@
This adjuster emphasizes only misclassified elements. However, the discontinuity of its corresponding adjusted loss function renders it unsuitable for gradient-based optimization. Therefore, a smoother, differentiable formulation is necessary.
3.3 Design of the Binarization-Aware Adjuster